Fermi National Laboratory
 

Accelerator Update for February 12, 2007 - February 14, 2007

 

Monday February 12

The day shift began with Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (14.2E10) and the Antiproton Source (Pbar) stack (156.7E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the colliding beams.

Operations established store 5224 at 8:24 AM with an initial luminosity of 252.5E30.

The Neutron Therapy Facility (NTF) reported at 11:52 AM that they had completed treatment on their last patient for the day.

MI experts put the MI-8 collimator into the beam

Pbar stopped stacking at 1:12 PM for studies. With Pbar off, Operations sent all of the beam to NuMI.

At 1:29 PM, Operations reported that the MI-8 line beam permit had been tripping off due to the affect the MI-8 collimators were having on the beam. It reset okay.

The Fire alarm technician reported at 3:03 PM that they were still working on the D0 VESDA alarm.

Operations reported at 3:56 PM that the Recycler had stashed antiprotons twice during the day shift.

Pbar resumed stacking at 4:48 PM.

At 5:19 PM, a Linac Klystron RF station (KRF6) tripped off on a LCW flow fault. Operators had the station recovered and back on within five minutes.

At 5:51 PM, the NuMI beam permit tripped due to a bad position on their target. It reset.

At 5:57 PM, Operations reported that MI had suffered a few aborts due to the MI-8 collimators. Operations reset the aborts.

At 8:41 PM, the LRF3 water skid tripped off. Operators investigated and found the control box without power. The operators got power back to the control box and the skid resumed normal operations, but then the skid developed an LCW leak. Operators turned off the skid and called in experts.

Linac experts reported at 10:42 PM that they were replacing a 24-volt power supply for the LRF3 water skid. They also explained that the LCW leak went away after the water skid was turned on and the water temperature rose.

Operations reported at midnight that the Recycler had only stashed once during the evening shift.

Tuesday February 13

The midnight shift began with Operations monitoring store 5224 (56.98E30) and the stash (210.6E10), and with Pbar not stacking (23.1E10) and NuMI off due to LRF3 repairs.

Linac experts completed their repairs on the LRF3 water skid at 12:35 AM. Operations resumed sending beam to Pbar and NuMI.

At 12:54 AM, MI tripped off for the first of many, many times during the midnight shift due to the MI-8 collimators. Operation reset MI each time.

At 3:36 AM, Operations reported that the H- Source was sparking and causing beam position problems. This problem continued all shift.

Two TeV experts began an end-of-store study at 6:08 AM.

Operations terminated store 5224 at 7:59 AM.

Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.

Operations prepared CDF and D0 for access at 8:15 AM.

At 9:16 AM, LRF4 tripped off for the first of three times in five minutes. It reset okay.

Cryo System technicians reported at 9:43 AM that they had replaced a wet engine flywheel for the TeV sector A2 cryo system.

At 10:15 AM, NuMI tripped off due to bad beam position on their target. It reset.

At 10:18 AM, the MI-8 line tripped off for the first of 21 times in 48 minutes due to the collimators. It reset each time.

CDF came out of access at 10:36 AM.

At 10:56 AM, NuMI tripped off due to a BPM failure. It reset okay.

At 10:58 AM, the H- Source sparked and caused NuMI to trip off. It reset.

At 11:17 AM, NuMI tripped off due to a quadrupole (QF112) power supply problem. Everything reset.

At 11:31 AM, NuMI tripped off due to a bad beam position on their target. It reset.

At 11:39 AM, Operations reported that their alarms program had been down for an hour. An expert investigated and had the program running within 18 minutes.

At 12:21 PM, Operations reported that the H- Source sparking was causing batch problems. An expert lowered the extractor voltage. (Two hours later, Operations reported that there had been no MI-8 collimator trips since this adjustment.)

Linac experts reported at 12:29 PM that the Klystron Buncher Coil power supply was about to fail. Since D0 was still in access, Operations told them to go ahead and begin repairs.

The Klystron Buncher work ended. Operations resumed sending beam to Pbar and NuMI at 1:53 PM.

D0 came out of access at 3:03 PM.

Operations put the TeV into a dry squeeze at 3:16 PM.

Operations put the TeV into shot setup at 3:50 PM.

Operations reported at 3:53 PM that the Recycler stashed twice during the day shift.

A Preaccelerator expert reported at 3:58 PM that an H- extractor tube was having intermittent problems. The experts said they would investigate further on Wednesday.

Operations established store 5226 at 6:16 PM with an initial luminosity of 284.5E30. This store set a New Luminosity Record!

At 11:10 PM, Operations reported that the TeV sector A1 wet engine was failing. Operators investigated and found that the engine was venting helium. Operations contacted the Cryo Coordinator.

At 11:49 PM, LRF4 tripped off. It reset.

Wednesday February 14

The midnight shift began with Operations monitoring store 5226 (118.99E30) and the stash (143.6E10), with Pbar stacking (23.1E10), with the TeV sector A1 wet engine failing, and with NuMI taking beam.

At 12:30 AM, LRF3 tripped off. It took about 50 minutes to reset the station.

At 1:31 AM, Operations reported problems loading ACNET programs. Operators investigated and then called an expert.

At 4:27 AM, Operations reported that a Pbar Debuncher quadrupole power supply (D:Q903) had tripped. Operations called in a power supply expert.

At 4:38 AM, store 5226was lost due to a quench at sector A1 due to the wet engine failing. Cryo system techs began changing out a leaking valve.

A power supply expert reported at 5:49 AM that D:Q903 was up and running.

At 5:57 AM, LRF5 tripped off. It reset.

Cryo system techs reported at 5:58 AM that they had completed their repairs on the TeV A1 wet engine.

At 6:04 AM, CHL reported that a warm helium compressor at TeV sector E2 had failed. CHL started up the E1 compressor.

At 6:17 AM, BRF6 tripped off. It reset.

Operations reported at 8 AM that the Recycler had stashed twice during the midnight shift.

The Plans for Wednesday

The plans for today are to wait for the TeV cryo to recover and then go into shot setup around noon.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/14/2007    email Fermilab